OpenStudy (anonymous):

Find f(-2) if f(x)=-x^3+2x^2+x-1

OpenStudy (anonymous):

just plug in the value -2 wherever you see x

OpenStudy (anonymous):

\[f(x) = -2^3 + (2*-2^2) -2 - 1\]

OpenStudy (anonymous):

just a mistake -x^3

OpenStudy (anonymous):

-13 ?

OpenStudy (anonymous):

f(-2)=-(-2)^3+2(-2)^2+(-2)-1 f(-2) = 2^3 + 2(-2)^2 -2 -1 f(-2) = 8 + 16 - 3 f(-2) = 24 - 3 f(-2) = ? subract 3 from 24 t get your answer

OpenStudy (anonymous):

let me know if something doesnt make sense

OpenStudy (anonymous):

13 was the correct answer

OpenStudy (anonymous):

yup 13

OpenStudy (anonymous):

true or false f(x)=3/x+7x^2 is a polynomial

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

do you know what polynomial means?

OpenStudy (anonymous):

yes an expression of more than two algebraic terms

OpenStudy (anonymous):

yes an expression of more than two algebraic terms

OpenStudy (anonymous):

polynomial means many terms "poly" meaning "many", "nomial" means term. Those terms can consist of, constant or variables or exponents but in this equation there is a problem look down in this picture that I attached

OpenStudy (anonymous):

ok then it is false

OpenStudy (anonymous):

can you give the reasoning?

OpenStudy (anonymous):

yes becasue the dividing in the problem

OpenStudy (anonymous):

correct 3/x cannot be considered as a term so it's no longer 'poly'nomial

OpenStudy (anonymous):

State the degree:rst+r^2s^2+r^3s^2 can you help me with this

OpenStudy (anonymous):

degree in this simply means the largest exponent of a particular variable \[r^2 + s^2 + r^3s^2\]

OpenStudy (anonymous):

ok its 5

OpenStudy (anonymous):

@what!? sorry I did notice the -x^3 later, after I posted, didn't feel like re-writing it.

OpenStudy (anonymous):

no you have to state for each variable separately

OpenStudy (anonymous):

is rst considered part of the equation in this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright. state the variables first then.